# ChatDBG: Augmenting Debugging with Large Language Models

### v3

@@ Abstract.
Debugging is a critical but challenging task for programmers. This paper proposes ChatDBG, an AI-powered debugging assistant. ChatDBG integrates large language models (LLMs) to significantly enhance the capabilities and user-friendliness of conventional debuggers. ChatDBG lets programmers engage in a collaborative dialogue with the debugger, allowing them to pose complex questions about program state, perform root cause analysis for crashes or assertion failures, and explore open-ended queries like ‘why is x null?’. To handle these queries, ChatDBG grants the LLM autonomy to take the wheel: it can act as an independent agent capable of querying and controlling the debugger to navigate through stacks and inspect program state. It then reports its findings and yields back control to the programmer. By leveraging the real-world knowledge embedded in LLMs, ChatDBG can diagnose issues identifiable only through the use of domain-specific reasoning. Our ChatDBG prototype integrates with standard debuggers including LLDB and GDB for native code and Pdb for Python. Our evaluation across a diverse set of code, including C/C++ code with known bugs and a suite of Python code including standalone scripts and Jupyter notebooks, demonstrates that ChatDBG can successfully analyze root causes, explain bugs, and generate accurate fixes for a wide range of real-world errors. For the Python programs, a single query led to an actionable bug fix 67% of the time; one additional follow-up query increased the success rate to 85%. ChatDBG has seen rapid uptake; it has already been downloaded more than 65,000 times.
@@
$c$s: Computing methodologies Artificial intelligenceccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Debuggers can be helpful, but finding and fixing software defects remains a deeply challenging and time-consuming task (Zeller, 2009a; Eisenstadt, 1993; Layman et al., 2013). Programmers must still reason about program behavior to ascertain what went wrong. They must formulate and test hypotheses about program execution, they must read and understand code they may have not written, and they must pore over potentially voluminous information. Such information includes lengthy executions, large amounts of program data, and many stack frames that potentially span multiple threads.
@@
A debugger integrated with ChatDBG continues to provide its full range of functionality but also lets programmers engage in debugging dialogs where they can ask high-level questions like ‘why is x null here?’ or ‘why isn’t this value what I expected?’. The question can be as simple as ‘why?’ if a program has crashed or failed an assertion. To answer such queries, ChatDBG orchestrates a conversation with an LLM. A key feature of ChatDBG is that it grants autonomy to the LLM to “take the wheel” and act as an independent agent (Wooldridge and Jennings, 1995; Gur et al., 2024) while answering the programmer’s queries. Specifically, the LLM issues “function calls” (OpenAI, 2023a) to run commands in the underlying debugger to investigate program state, execute code, or obtain source code. The results of those calls are sent back to the LLM for use while constructing its response. After answering a query, control is returned to the programmer, who may then enter additional commands or chat messages.
Our prototype of ChatDBG integrates into three widely used debuggers: GDB, LLDB, and Pdb. Our evaluation presents a range of case studies demonstrating that ChatDBG improves significantly on existing debuggers. On a suite of unpublished Python scripts and Jupyter notebooks written by undergraduate students, one or two queries is sufficient for ChatDBG to properly diagnose and fix defects 87% of the time, typically at a cost well under \$0.20 USD. ChatDBG is also effective at identifying causes and providing fixes for a range of real-world bugs in C/C++ code.
This paper makes the following contributions.
@@
This paper is organized as follows. Section 2 illustrates the use of ChatDBG to debug a program. Section 3 describes key related work. Section 4 describes ChatDBG’s implementation. Section 5 presents our evaluation. Section 6 outlines directions for future work, and Section 7 concludes.
@@ 2. Overview
ChatDBG identifies the root cause: proportion_blue fails to compute and return the desired statistic, and it provides a corrected version of proportion_blue. When ChatDBG cannot identify the root cause, it suggests further debugging steps and control is returned to the user, who may continue the chat, issue further debugger commands, or both. Figure 3 illustrates this scenario, where a version of bootstrap.py with the corrected proportion_blue function fails the assertion on line 21. The user asks why the mean of stats is not close to 0.7, and ChatDBG’s first response suggests examining whether 0.7 is the appropriate expected value. The user then computes the proportion of blue marbles with a debugger command and tells ChatDBG that 0.7 is the correct value. In response, ChatDBG points to the low number of trials (5) as the issue. The LLM drew this correct conclusion without seeing any discussion of trial size or variance in any program state, code, or documentation encountered during the chat. A powerful aspect of ChatDBG is its ability to exploit real-world knowledge in its analyses (here, the fact that bootstrapping depends on large numbers of resamples) without specific instruction or user intervention.
@@ 4.6. Navigating the Code
ChatDBG augments Pdb with the info command, which prints the docstring for any function, class, field, method, or package. It additionally prints the source for any user-defined code. The info requests in Figure 2 demonstrate these two cases for proportion_blue and bootstrap_statistic, respectively. The command is implemented via the standard inspect and pydoc libraries.
@@ 4.7. Slices for Interactive Python
ChatDBG supports debugging interactive IPython sessions and Jupyter notebooks. Interactive sessions lead to many individual code cells that are each evaluated separately. Cells may be evaluated out-of-order, override definitions from earlier cells, and communicate values to other cells through top-level global variables. Others have noted the challenges of reasoning about program behavior in this context (Head et al., 2019; Shankar et al., 2022). ChatDBG provides an additional slice debugging command to facilitate that reasoning. The slice command computes the backwards slice for any variable used in the current cell that was defined in previously-executed cells. It returns the code for cells in that slice. Suppose the code from bootstrap.py in Figure 1 were written in four notebook cells:
@@ 4.8. Security and Risks
s1 123 semantic Assertion Error Incorrect drop and rename operations leading to bad data in table
@@
s11 181 crash Value Error Creates a sample without replacement that is larger than the input
@@
s14 65 crash Index Error Incorrect computation of a random sample under null hypothesis
@@ 5.1. Python
We applied ChatDBG to all bugs in a collection of student labs from two introductory computer science courses; see Table 3. Bugs c1–c8 are in non-interactive scripts from a programming class that perform various file reading and text processing tasks. Bugs s1–s14 are in Jupyter notebooks (Team, 2015) from a data science class that manipulate, visualize, and compute over arrays and tables. Some bugs were apparent to the programs’ authors. Others were identified during autograding.
@@
ChatDBG used the gpt-4-1106-preview model for these experiments. Under +Targeted Question, the first prompt and response led to, on average, a chat of about 10,000 tokens (7,500 words), a cost of about \$$0.120.12$ USD under OpenAI’s current pricing model (OpenAI, [n. d.]), and a completion time of about 25 seconds. Subsequent steps in extended debugging dialogs incurred comparable costs. Time was highly variable and dominated by the performance of OpenAI’s service. These characteristics will be different for other platforms and models and, given current trends, we expect significant reductions in both time and cost as models improve.
@@
Failures were generally due to the LLM not always recognizing or discovering key aspects of a program’s behavior. We observe that in some cases, enabling additional features in ChatDBG decreases its success rate. We attribute this result to the fact that longer and more complex prompts can occasionally degrade the effectiveness of LLMs (Levy et al., 2024). In general, the further the distance between the root cause of a bug and observable effect, the more challenging it was for ChatDBG (and people (Zeller, 2009b, p. 243)) to find it. In some cases, it was on the right track but did not converge on an actionable fix. In others, it suggested changes that would introduce other bugs. It also occasionally made mistakes, such as conflating proportions and percentages or failing to handle unusual corner cases. All of these could be mitigated by feedback from the user in subsequent follow-ups.
5.2. C and C++
@@ 7. Conclusion
ChatDBG is available on GitHub at github.com/plasma-umass/ChatDBG.

### v5

@@ Abstract.
Debugging is a critical but challenging task for programmers. This paper proposes ChatDBG, an AI-powered debugging assistant. ChatDBG integrates large language models (LLMs) to significantly enhance the capabilities and user-friendliness of conventional debuggers. ChatDBG lets programmers engage in a collaborative dialogue with the debugger, allowing them to pose complex questions about program state, perform root cause analysis for crashes or assertion failures, and explore open-ended queries like ‘why is x null?’. To handle these queries, ChatDBG grants the LLM autonomy to take the wheel: it can act as an independent agent capable of querying and controlling the debugger to navigate through stacks and inspect program state. It then reports its findings and yields back control to the programmer. By leveraging the real-world knowledge embedded in LLMs, ChatDBG can diagnose issues identifiable only through the use of domain-specific reasoning. Our ChatDBG prototype integrates with standard debuggers including LLDB and GDB for native code and Pdb for Python. Our evaluation across a diverse set of code, including C/C++ code with known bugs and a suite of Python code including standalone scripts and Jupyter notebooks, demonstrates that ChatDBG can successfully analyze root causes, explain bugs, and generate accurate fixes for a wide range of real-world errors. For the Python programs, a single query led to an actionable bug fix 67% of the time; one additional follow-up query increased the success rate to 85%. ChatDBG has seen rapid uptake; it has already been downloaded more than 75,000 times.
@@
$c$s: Computing methodologies Artificial intelligenceccs: Software and its engineering Software testing and debuggingcopyright: ccdoi: 10.1145/3729355journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE085publicationmonth: 7
@@ 1. Introduction
Debuggers can be helpful, but finding and fixing software defects remains a deeply challenging and time-consuming task (Zeller, 2009; Eisenstadt, 1993; Layman et al., 2013). Programmers must still reason about program behavior to ascertain what went wrong. They must formulate and test hypotheses about program execution, they must read and understand code they may have not written, and they must pore over potentially voluminous information. Such information includes lengthy executions, large amounts of program data, and many stack frames that potentially span multiple threads.
@@
A debugger integrated with ChatDBG continues to provide its full range of functionality but also lets programmers engage in debugging dialogs where they can ask high-level questions like ‘why is x null here?’ or ‘why isn’t this value what I expected?’. The question can be as simple as ‘why?’ if a program has crashed or failed an assertion. To answer such queries, ChatDBG orchestrates a conversation with an LLM. A key feature of ChatDBG is that it grants autonomy to the LLM to “take the wheel” and act as an independent agent (Wooldridge and Jennings, 1995; Gur et al., 2024) while answering the programmer’s queries. Specifically, the LLM issues “function calls” (OpenAI, 2023a) to run commands in the underlying debugger to investigate program state, execute code, or obtain source code. The results of those calls are sent back to the LLM to use in constructing its response. After answering a query, control is returned to the programmer, who may then enter additional commands or chat messages.
Our prototype of ChatDBG integrates into three widely used debuggers: GDB, LLDB, and Pdb. Our evaluation presents a range of case studies demonstrating that ChatDBG improves significantly on existing debuggers. On a suite of unpublished Python scripts and Jupyter notebooks written by undergraduate students, one or two queries is sufficient for ChatDBG to properly diagnose and fix defects 85% of the time, typically at a cost well under \$0.20 USD. ChatDBG is also effective at identifying causes and providing fixes for a range of real-world bugs in C/C++ code.
This paper makes the following contributions:
@@ 2. Overview
[…]
Text snippet of a ChatDBG debugging session.
ChatDBG identifies and corrects the root cause: proportion_blue incorrectly computes the desired statistic. When ChatDBG cannot identify the root cause, it suggests further debugging steps and control is returned to the user, who may continue the chat, issue further debugger commands, or both. Figure 3 illustrates this scenario, where a version of bootstrap.py with the corrected proportion_blue function fails the assertion on line 21.
The user asks why the mean of stats is not close to 0.7, and ChatDBG’s initial response suggests examining whether 0.7 is the appropriate expected value. To verify this, the user then computes the proportion of blue marbles with a debugger command and tells ChatDBG that 0.7 is indeed the correct value. In its follow-up, ChatDBG points to the low number of trials (five) as the issue. The LLM drew this correct conclusion without seeing any discussion of trial size or variance in any program state, code, or documentation encountered during the chat. A powerful aspect of ChatDBG is its ability to exploit real-world knowledge in its analyses (here, the fact that bootstrapping depends on large numbers of resamples) without specific instruction or user intervention.
@@ 4.2. Debugging a Target Program
[…]
A graph showing the architecture of ChatDBG, along with a short algorithm summary. Users may directly issue standard debugger commands, that are handled by the existing debugger. They may also use free-form text, that will be interpreted by ChatDBG. ChatDBG will act as an agent, and facilitate conversation between the LLM and the underlying debugger. The LLM will issue commands to retrieve context about source code or runtime state before issuing its final response and/or recommendation.
@@ 4.6. Navigating the Code
ChatDBG augments Pdb with the info command, which prints the docstring for any function, class, field, method, or package. It additionally prints the source code for any user-defined function. The info requests in Figure 2 demonstrate these two cases for proportion_blue and bootstrap_statistic, respectively. The command is implemented via the standard inspect and pydoc Python libraries.
@@ 4.7. Slices for Interactive Python
ChatDBG supports debugging interactive IPython sessions and Jupyter notebooks. Interactive sessions lead to many individual code cells that are each evaluated separately. Cells may be evaluated out-of-order, override definitions from earlier cells, and communicate values to other cells through top-level global variables. Others have noted the challenges of reasoning about program behavior in this context (Head et al., 2019; Shankar et al., 2022). ChatDBG provides an additional slice debugging command to facilitate that reasoning. The slice command computes the backwards slice for any variable used in the current cell that was defined in previously-executed cells. It returns the code for cells in that slice. Suppose the code from bootstrap.py in Figure 1 were written in four notebook cells as shown below:
@@ 4.8. Security and Risks
s1 123 semantic Assertion Error Incorrect drop and rename operations leading to bad data
@@
s11 181 crash Value Error Creates a sample without replacement larger than the input
@@
s14 65 crash Index Error Incorrect computation of random sample under null hypothesis
@@ 5.1. Python
We applied ChatDBG to all of the bugs in a collection of student labs from two introductory computer science courses; see Table 3. Bugs c1–c8 are in non-interactive scripts from a programming class that perform various file reading and text processing tasks. Bugs s1–s14 are in Jupyter notebooks (Team, 2015) from a data science class that manipulate, visualize, and compute over arrays and tables. Some bugs were apparent to the programs’ authors. Others were identified during autograding.
@@
ChatDBG used the gpt-4-1106-preview model for these experiments. Under +Targeted Question, the first prompt and response led to, on average, a chat of about 10,000 tokens (7,500 words), a cost of about \$$0.120.12$ USD under OpenAI’s current pricing model (OpenAI, 2024), and a completion time of about 25 seconds. Subsequent steps in extended debugging dialogs incurred comparable costs. Time was highly variable and dominated by the performance of OpenAI’s service. These characteristics will be different for other platforms and models and, given current trends, we expect significant reductions in both time and cost as models improve.
Table 4. Configurations used in the Python experiments (§5.1).
@@
[…]
Chart showing the overall success rate of ChatDBG under each configuration. Each feature addition improves the success rate, culminating at 85% using the fully-featured ChatDBG.
@@
[…]
A detailed version of Figure 6, showing the success rate for each program evaluated under each configuration. While there is some variation, the ChatDBG generally improves with additional features.
Failures were generally due to the LLM not always recognizing or discovering key aspects of a program’s behavior. We observe that in some cases, enabling additional features in ChatDBG decreases its success rate. We attribute this result to the fact that longer and more complex prompts can occasionally degrade the effectiveness of LLMs (Levy et al., 2024). In general, the further the distance between the root cause of a bug and observable effect, the more challenging it was for ChatDBG (and people (Zeller, 2009, p. 243)) to find it. In some cases, it was on the right track but did not converge on an actionable fix. In others, it suggested changes that would introduce other bugs. It also occasionally made mistakes, such as conflating proportions and percentages or failing to handle unusual corner cases. All of these could be mitigated by feedback from the user in subsequent follow-ups.
5.2. C and C++
@@
[…]
Chart showing the success rate of ChatDBG at fixing the proximate or root cause in C/C++ programs. Proximate causes are identified and fixed most of the time, while success finding and fixing root cause varies among programs.
@@ 7. Conclusion
Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grant No. 2243636. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ Acknowledgements.
ChatDBG is available on GitHub at github.com/plasma-umass/ChatDBG. An archived version is also available on Zenodo (Levin et al., 2025).